# Metric theory of inhomogeneous Diophantine approximations with a fixed matrix

### v2

@@ 4.2 Criteria of badly and very well approximability
(Theorem 2.1 from [17]) Matrix $\Theta$ is badly approximable if and only if there exists $C>0$, such that
@@ 4.3 On a result by Bugeaud and Laurent
$\omega(\Theta,\boldsymbol{\eta})\geq\frac{1}{\hat{\omega}(\Theta^{\top})}\,\,\,\,\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,\,\,\,\,\hat{\omega}(\Theta,\boldsymbol{\eta})\geq\frac{1}{\omega(\Theta^{\top})}.$ (21)
Moreover, let $\mathcal{A}$ be a non-$\Xi$-exceptional affine subspace of $\mathbb{R}^{n}$ for some $0<\Xi<1$. Then, for Lebesgue-almost all $\boldsymbol{\eta}\in\mathcal{A}$, the equality in (21) holds.
@@ 4.4 On a result by Kleinbock and Wadleigh
One can show that our Theorem 1 together with the Khintchine–Groshev theorem provide an immediate proof of Theorem G, and vice versa: Khintchine–Groshev theorem follows from Theorem G via Theorem 1. The detailed proof of this equivalence in a more general case of approximations with general weight functions can be found in [43]; here we will just sketch the idea of the argument.
First, one can check that it is enough to show Theorem G for strictly decreasing continuous functions $g$: the general is an easy corollary. From now on, assume $g$ is strictly decreasing and continuous. Let $f$ and $g$ will be connected by (2). In Theorem G, we are only interested in the values of $g$ at integer points, so we can assume that $g$ (and thus $f$ as well) is a differentiable function.
Let us notice that by the variable change $t=\frac{1}{g(z)}=f^{-1}\left(\frac{1}{z}\right)$ the integrals
@@
either both converge or both diverge. Notice also that the convergence (or divergence) of these integrals is invariant under multiplication of $f$ or $g$ by a constant and under a linear variable change. Let us fix $\varepsilon<\frac{1}{4}$.
@@
  • Suppose $\sum\limits_{j=1}^{\infty}\frac{1}{j^{2}g(j)}=\infty$; then, by Khintchine–Groshev theorem, for almost every $\Theta\in{\bf M}_{n,m}$, its transposed matrix $\Theta^{\top}\in M_{m,n}$ belongs to ${\bf W}_{m,n}\left[\delta f(KT)\right]$ for any constants $\delta,K>0$. Choosing proper values of $\delta$ and $K$ and applying Theorem 1, for almost every $\Theta\in{\bf M}_{n,m}$ one has
    $m\left(\widehat{\bf D}_{n,m}^{\Theta}\left[g\right]\right)=0.$
    The set $\widehat{\bf D}_{n,m}\left[g\right]$ is Borel and thus measurable. Fubini’s theorem applied to the characteristic function of this set shows that
    $m\left(\widehat{\bf D}_{n,m}\left[g\right]\right)=0.$
  • Suppose $\sum\limits_{j=1}^{\infty}\frac{1}{j^{2}g(j)}<\infty$; then, by Khintchine–Groshev theorem, for any $C>0$ we have
    $m\left({\bf W}_{m,n}\left[Cf(T)\right]\right)=0,$
    which means that for almost every $\Theta\in M_{n,m}$
    $\Theta^{\top}\in{\bf BA}_{n,m}\left[f(T);C\right].$
    By Theorem A then, for almost every $\Theta\in M_{n,m}$
    $\widehat{\bf D}_{n,m}^{\Theta}\left[g\left(\frac{C}{\kappa^{m}}T\right)\right]=\mathbb{R}^{n}.$
    We take $C=\kappa^{m}$ to conclude that $\widehat{\bf D}_{n,m}^{\Theta}\left[g\right]=\mathbb{R}^{n}$. It remains to apply Fubini’s Theorem again to complete the proof.
@@ 5.1 Irrationality measure function
Let us notice that we can equivalently define the Diophantine exponents using the irrationality measure function. It is easy to see that
@@ 5.2 Best approximations: notation and general properties
If $\Theta$ is not trivially singular, there exists the sequence
@@
Let values $A$ and $B$ be defined in (17) Then for any $\Theta$ which is not trivially singular and for any $\nu$ one has
  1. A.
@@
The first statement can be found in [7]. The second statement is quite similar but we do not know where it is documented. So we give a proof here. This proof generalizes the argument from [39]. Indeed, let us consider boxes
@@
By the definition of the best approximation there are no non-zero integer points strictly inside $\Pi_{1}$. The set $\Pi_{1}\setminus\Pi_{2}$ can be covered by exactly $B_{1}=2^{2n}-2^{n}$ shifts of the box
@@
$\boldsymbol{\xi}_{\nu+j}=\Theta{\bf p}_{\nu+j}-{\bf{a}}_{\nu+j},\,\,\,0\leqslant j\leqslant B.$ (22)
If at leas one of the remainders (22) belongs to $\Pi_{2}$, everything is proven. So we may suppose that all the remainders (22) belong to $\Pi_{1}\setminus\Pi_{2}$. By pigeonhole principle, there exist a shift $\Pi_{0}^{*}$ of $\Pi_{0}$ which contains at least $\frac{B}{B_{1}}+1=2^{m-1}+1$ different remainder vectors
@@
of the form (22). For all these vectors we have
@@
Applying pigeonhole principle one more time for the collection of vectors
$\pm p_{j_{k}},\,\,\,1\leqslant k\leqslant 2^{m-1}+1,$ (23)
we see that there are at least two integer vectors of the form (23), say ${\bf p}_{\nu+j^{\prime}}$ and ${\bf p}_{\nu+j^{\prime\prime}}$ which belong to the same shift of $\mathcal{P}_{0}$ and so
@@
By the definition of the best approximation this means that $\psi_{\Theta}(P_{\nu+B})/2\leqslant\psi_{\Theta}(P_{\nu})/2$.
@@
Throughout the paper, we will keep this notation ${\bf q}_{\nu}$ for the best approximations vectors of the pair $(\Theta,\boldsymbol{\eta})$. We will also consistently use the notation ${\bf p}_{\nu}$ for the best approximations vectors of a matrix $\Theta$, given it is not trivially singular, and $P_{\nu}=|{\bf p}_{\nu}|$. We will use the notation ${\bf y}_{\nu}$ and $Y_{\nu}=|{\bf y}_{\nu}|$ analogously for the best approximation vectors of the transposed matrix $\Theta^{\top}$ and their magnitudes.
@@ 5.3 Asymptotic directions
In Theorem 2 and Theorem 4, the condition on $\mathcal{A}$ depends on the set of all the asymptotic directions of the best approximation vectors of $\Theta^{\top}$. In this subsection, we give the definitions and brief overview of the subject.
Let $\Theta^{\top}$ be a non trivially singular matrix, and ${\bf y}_{\nu}$ its sequence of the best approximation vectors. Let $S^{n-1}$ be the n1n-1-dimensional sphere in $\mathbb{R}^{n}$ centered at zero. We say $\zeta$ is an asymptotic direction for the best approximations sequence ${\bf y}_{\nu}$ if $\zeta$ is a limit point of the set $\left\{\pm\frac{{\bf y}_{\nu}}{|{\bf y}_{\nu}|}:\,\,\nu\in\mathbb{N}\right\}$. We will denote the set of all the asymptotic direction by $\Omega$, and will refer to it shortly just as "the set of asymptotic directions".
@@
The sets of asymptotic directions were studied before; see [19] for the results concerning the asymptotic directions sets in the case of linear forms. It is easy to see that any set of asymptotic directions is closed and centrally symmetric. It is shown in [19] that in the special case $m=1$ there are no other restrictions:
@@
(Theorem 4 in [19]) Let $\Omega\subseteq S^{n-1}$ be a closed centrally symmetric set. Then, the set of $1\times n$ matrices $\Theta^{\top}$ for which $\Omega$ is the set of asymptotic directions has cardinality continuum.
In some special cases, we can construct the linear forms with the prescribed set of the asymptotic directions and in addition control its approximation properties. Namely, Theorem 2 from [20] guarantees that for any $\gamma$ large enough there exists such a $\Theta^{\top}\in M_{1,n}$ that $\omega(\Theta^{\top})=\gamma$ and the set $\Omega$ of the asymptotic directions of $\Theta^{\top}$ consists of two pairs of centrally symmetric points. It is easy to see from the proof that we can control the uniform exponent of $\Theta^{\top}$ or precise function $f$ for which $\Theta^{\top}$ is $f$-approximable/$f$-Dirichlet/etc. instead. This provides us with the examples for which Theorem 2 and Theorem 4 hold for almost every $\mathcal{A}$.
@@ 6.1 Exceptional sets
In this section, we introduce some useful language and prove simple but useful statements.
First, assume $\Theta^{\top}$ is not trivially singular. Let us call a point ${\bf v}\in\mathbb{R}^{n}$ on the unit sphere $\Xi$-exceptional (for the matrix $\Theta^{\top}$) if
$\text{for any}\,\,\delta>0\,\,\text{ there exist infinitely many }\,\,\ \nu\in\mathbb{N}\,\,\text{for which}\,\,\widehat{{\bf y}_{\nu},{\bf v}}<\frac{\delta}{|{\bf y}_{\nu}|^{\Xi}}.$ (24)
@@
We will treat the case of trivially singular $\Theta^{\top}$ separately. Suppose ${\bf 0\neq z}\in\mathbb{Z}^{n}$ is such a vector that $\Theta^{\top}{\bf z}={\bf 0}$; we will say that the exceptional subspaces are the subspaces of the form
$\{\boldsymbol{\eta}\in\mathbb{Z}^{n}:\,\,\,\boldsymbol{\eta}\cdot{\bf z}=a\}\,\,\,\,\,\,\,\,\,\,\,\,\text{for ${\bf a\in\mathbb{Z}}$}$ (25)
or the ones that are contained in the subspaces of the form (25). For convenience, we will use a convention that exceptional sets are $\Xi$-exceptional for any $\Xi$.
We prove our main theorems on subspaces which are not exceptional. Here, we aim to show that the sets of $\Xi$-exceptional subspaces are very small.
@@
The set of the $\Xi$-exceptional points on the unit sphere has zero Hausdorff dimension for any $\Xi>0$.
@@
Let us denote the set of the exceptional points by $E$. Then,
@@
Fix $s>0$, and let $C(s):=\sum\limits_{\nu=1}^{\infty}\left(\frac{\delta}{|{\bf y}_{\nu}|^{\Xi}}\right)^{s}$. For any $\delta>0$,
@@
where the limit being zero is provided by the first statement of Proposition 1. This shows that $s$-dimensional Hausdorff measure of $E$ is zero and completes the proof. ∎
@@
Let $e\in E$; then the set of all the $r$-dimensional subspaces containing $e$ can be identified with $Gr(l,r)$. Let us consider the set $P_{E}=\{(e,V):\,\,\,e\in E,\,e\in V,\,V\in X_{\{e\}}\}$: it can be identified with $E\times Gr(l,r)$. Recall that the Hausdorff and upper box dimensions for the set $Gr(l,r)$ coincide; therefore, $\dim_{H}P_{E}=\dim_{H}Gr(l,r)+\dim_{H}E=r(l-r)$. The set $X_{E}$ is the projection of $P_{E}$ on the second coordinate, so $\dim_{H}X_{E}\leq\dim_{H}P_{E}=r(l-r)$.
From the other hand, let us notice that the set $X_{\{e\}}$ of all $r$-dimensional subspaces containing a single point $e$ can be identified with $Gr(l,r)$ and thus has dimension r(lr)r(l-r). But $E$ is nonempty, thus $X_{E}$ contains a set of the form $X_{\{e\}}$ and thus $\dim_{H}X_{E}\geq\dim_{H}X_{\{e\}}=r(l-r)$, which completes the proof. ∎
@@
Let $X_{E}$ be the set of $k$-dimensional linear subspaces of $\mathbb{R}^{n}$ intersecting with the set $E$ of $\Xi$-exceptional points for $\Theta^{\top}$. Then, $\dim_{H}X_{E}=(k-1)(n-k)$, if not empty.
Corollary 3 shows that the linear subspaces $\mathcal{L}$ corresponding to $\Xi$-exceptional $k$-dimensional affine subspaces $\mathcal{A}$ form a $\leq(k-1)(n-k)$-dimensional subspace on the grassmanian $Gr(n,k)$ for any $\Xi>0$. In particular, the set of $\Xi$-exceptional $k$-dimensional affine subspaces has measure zero.

### v3

@@ 4.2 Criteria of bad and very good approximability
(Theorem 2.1 from [19]) A matrix $\Theta$ is badly approximable if and only if there exists $C>0$ such that
@@ 4.3 On a result by Bugeaud and Laurent
$\omega(\Theta,\boldsymbol{\eta})\geq\frac{1}{\hat{\omega}(\Theta^{\top})}\,\,\,\,\,\,\,\,\,\,\,\,\text{and}\,\,\,\,\,\,\,\,\,\,\,\,\hat{\omega}(\Theta,\boldsymbol{\eta})\geq\frac{1}{\omega(\Theta^{\top})}.$ (4.4)
Moreover, let $\mathcal{A}$ be a non-$\Xi$-exceptional affine subspace of $\mathbb{R}^{n}$ for some $0<\Xi<1$. Then, for Lebesgue-almost all $\boldsymbol{\eta}\in\mathcal{A}$, equality in (4.4) holds.
@@ 4.4 On a result by Kleinbock and Wadleigh
One can show that our Theorem 1 together with the Khintchine–Groshev theorem provide an immediate proof of Theorem G, and vice versa: Khintchine–Groshev theorem follows from Theorem G via Theorem 1. The detailed proof of this equivalence in a more general case of approximations with general weight functions can be found in [45]; here we will just sketch an idea of the argument.
First, one can check that it is enough to show Theorem G for strictly decreasing continuous functions $g$: the general case is an easy corollary. From now on, assume $g$ is strictly decreasing and continuous. Let $f$ and $g$ be related via (1.2). In Theorem G, we are only interested in the values of $g$ at integer points, so we can assume that $g$ (and thus $f$ as well) is a differentiable function.
Let us notice that by the change of variables $t=\frac{1}{g(z)}=f^{-1}\left(\frac{1}{z}\right)$ the integrals
@@
either both converge or both diverge. Notice also that the convergence (or divergence) of these integrals is invariant under multiplication of $f$ or $g$ by a constant and under a linear change of variables. Let us fix $\varepsilon<\frac{1}{4}$.
@@
  • Suppose $\sum\limits_{j=1}^{\infty}\frac{1}{j^{2}g(j)}=\infty$; then, by Khintchine–Groshev theorem, for almost every $\Theta\in{\bf M}_{n,m}$, its transposed matrix $\Theta^{\top}\in M_{m,n}$ belongs to ${\bf W}_{m,n}\left[\delta f(KT)\right]$ for any constants $\delta,K>0$. Applying Theorem 1 with a right choice of $\delta$ and $K$, one can show that $m\left(\widehat{\bf D}_{n,m}^{\Theta}\left[g\right]\right)=0$ for almost every $\Theta\in{\bf M}_{n,m}$. The set $\widehat{\bf D}_{n,m}\left[g\right]$ is Borel and thus measurable. Fubini’s theorem applied to the characteristic function of this set shows that $m\left(\widehat{\bf D}_{n,m}\left[g\right]\right)=0.$
  • Suppose $\sum\limits_{j=1}^{\infty}\frac{1}{j^{2}g(j)}<\infty$; then, by Khintchine–Groshev theorem, for any $C>0$ we have $m\left({\bf W}_{m,n}\left[Cf(T)\right]\right)=0,$ which means that $\Theta^{\top}\in{\bf BA}_{n,m}\left[f(T);C\right]$ for almost every $\Theta\in M_{n,m}$. By Theorem A then, for almost every $\Theta\in M_{n,m}$ we have that $\widehat{\bf D}_{n,m}^{\Theta}\left[g\left(\frac{C}{\kappa^{m}}T\right)\right]=\mathbb{R}^{n}.$ We take $C=\kappa^{m}$ to conclude that $\widehat{\bf D}_{n,m}^{\Theta}\left[g\right]=\mathbb{R}^{n}$. It remains to apply Fubini’s Theorem again to complete the proof.
@@ 5.1 Irrationality measure function
Let us also notice that we can equivalently define Diophantine exponents using the irrationality measure function. It is easy to see that
@@ 5.2 Best approximations: notation and general properties
If $\Theta$ is not trivially singular, there exists a sequence
@@
Let the values $A$ and $B$ be defined in (3.5) Then for any $\Theta$ which is not trivially singular and for any $\nu$ one has
  1. (a)
@@
The first statement can be found in [7]. The second statement is quite similar, but we do not know where it is documented. So we give a proof here. This proof generalizes the argument from [41]. Indeed, let us consider boxes
@@
By the definition of the best approximations there are no non-zero integer points strictly inside $\Pi_{1}$. The set $\Pi_{1}\setminus\Pi_{2}$ can be covered by exactly $B_{1}=2^{2n}-2^{n}$ shifts of the box
@@
$\boldsymbol{\xi}_{\nu+j}=\Theta{\bf p}_{\nu+j}-{\bf{a}}_{\nu+j},\,\,\,0\leqslant j\leqslant B.$ (5.1)
If at least one of the remainders (5.1) belongs to $\Pi_{2}$, everything is proven. So we may suppose that all the remainders (5.1) belong to $\Pi_{1}\setminus\Pi_{2}$. By the pigeonhole principle, there exists a shift $\Pi_{0}^{*}$ of $\Pi_{0}$ which contains at least $\frac{B}{B_{1}}+1=2^{m-1}+1$ different remainder vectors
@@
of the form (5.1). For all these vectors we have
@@
Applying the pigeonhole principle one more time to the collection of vectors
$\pm p_{j_{k}},\,\,\,1\leqslant k\leqslant 2^{m-1}+1,$ (5.2)
we see that there are at least two integer vectors of the form (5.2), say ${\bf p}_{\nu+j^{\prime}}$ and ${\bf p}_{\nu+j^{\prime\prime}}$, which belong to the same shift of $\mathcal{P}_{0}$, and thus
@@
By the definition of the best approximations this means that $\psi_{\Theta}(P_{\nu+B})/2\leqslant\psi_{\Theta}(P_{\nu})/2$.
@@
Throughout the paper, we will keep this notation ${\bf q}_{\nu}$ for the best approximation vectors of the pair $(\Theta,\boldsymbol{\eta})$. We will also consistently use the notation ${\bf p}_{\nu}$ for the best approximation vectors of a matrix $\Theta$, given it is not trivially singular, and will let $P_{\nu}=|{\bf p}_{\nu}|$. We will use the notation ${\bf y}_{\nu}$ and $Y_{\nu}=|{\bf y}_{\nu}|$ analogously for the best approximation vectors of the transposed matrix $\Theta^{\top}$ and their magnitudes.
@@ 5.3 Asymptotic directions
In Theorem 2 and Theorem 4, the condition on an affine subspace $\mathcal{A}$ of $\mathbb{R}^{n}$ depends on the set of all the asymptotic directions of the best approximation vectors of $\Theta^{\top}$. In this subsection, we give all the definitions and brief overview of the subject.
Let $\Theta^{\top}$ be a non trivially singular matrix, and ${\bf y}_{\nu}$ its sequence of the best approximation vectors. Let $S^{n-1}$ be the (n1)(n-1)-dimensional sphere in $\mathbb{R}^{n}$ centered at zero. We say $\zeta$ is an asymptotic direction for the best approximation sequence ${\bf y}_{\nu}$ if $\zeta$ is a limit point of the set $\left\{\pm\frac{{\bf y}_{\nu}}{|{\bf y}_{\nu}|}:\,\,\nu\in\mathbb{N}\right\}$. We will denote the set of all the asymptotic directions by $\Omega$, and will refer to it briefly just as "the set of asymptotic directions".
@@
The sets of asymptotic directions were studied before; see [21] for the results concerning the asymptotic direction sets in the case of linear forms. It is easy to see that any set of asymptotic directions is closed and centrally symmetric. It is shown in [21] that in the special case $m=1$ there are no other restrictions:
@@
(Theorem 4 in [21]) Let $\Omega\subseteq S^{n-1}$ be a closed centrally symmetric set. Then, the set of $1\times n$ matrices $\Theta^{\top}$ for which $\Omega$ is the set of asymptotic directions has the cardinality of continuum.
In some special cases, we can construct linear forms with a prescribed set of asymptotic directions and in addition control their approximation properties. Namely, Theorem 2 from [22] guarantees that for any large enough $\gamma$ there exists a $\Theta^{\top}\in M_{1,n}$ such that $\omega(\Theta^{\top})=\gamma$ and the set $\Omega$ of asymptotic directions of $\Theta^{\top}$ consists of two pairs of centrally symmetric points. It is easy to see from the proof that we can control the uniform exponent of $\Theta^{\top}$ or, more generally, a particular function $f$ for which $\Theta^{\top}$ is $f$-approximable/$f$-Dirichlet/etc. instead. This provides us with examples for which Theorem 2 and Theorem 4 hold for almost every $\mathcal{A}$.
@@ 6.1 Exceptional sets
In this section, we introduce some useful language and prove simple but useful statements. We use the notation $\measuredangle({\bf u,v})$ for the angle between two vectors ${\bf u,v}\in\mathbb{R}^{n}$.
First, assume $\Theta^{\top}$ is not trivially singular, and let $\Xi>0$. Let us call a point ${\bf v}\in\mathbb{R}^{n}$ on the unit sphere $\Xi$-exceptional (for the matrix $\Theta^{\top}$) if
$\text{for any}\,\,\delta>0\,\,\text{ there exist infinitely many }\,\,\ \nu\in\mathbb{N}\,\,\text{for which}\,\,\measuredangle({{\bf y}_{\nu},{\bf v}})<\frac{\delta}{|{\bf y}_{\nu}|^{\Xi}}.$ (6.1)
@@
We will treat the case of trivially singular $\Theta^{\top}$ separately. Suppose ${\bf 0\neq z}\in\mathbb{Z}^{n}$ is such that $\Theta^{\top}{\bf z}={\bf 0}$; we will say that the exceptional subspaces are the subspaces of the form
$\{\boldsymbol{\eta}\in\mathbb{Z}^{n}:\,\,\,\boldsymbol{\eta}\cdot{\bf z}=a\}\,\,\,\,\,\,\,\,\,\,\,\,\text{for ${\bf a\in\mathbb{Z}}$}$ (6.2)
or the ones that are contained in subspaces of the form (6.2). For convenience, we will use a convention that exceptional sets are $\Xi$-exceptional for any $\Xi$.
We prove our main theorems for subspaces that are not exceptional. Here, we aim to show that the sets of $\Xi$-exceptional subspaces are very small.
@@
The set of $\Xi$-exceptional points on the unit sphere has zero Hausdorff dimension for any $\Xi>0$.
@@
Let us denote the set of $\Xi$-exceptional points by $E$. Then,
@@
Fix $s>0$. Let $\delta>0$ and let $C(s):=\sum\limits_{\nu=1}^{\infty}\left(\frac{\delta}{|{\bf y}_{\nu}|^{\Xi}}\right)^{s}$. Then,
@@
where the limit being zero is provided by the first statement of Proposition 1. This shows that the $s$-dimensional Hausdorff measure of $E$ is zero and completes the proof. ∎
@@
Let $e\in E$; then the set of all the $r$-dimensional subspaces containing $e$ can be identified with $Gr(l,r)$. Let us consider the set $P_{E}=\{(e,V):\,\,\,e\in E,\,e\in V,\,V\in X_{\{e\}}\}$: it can be identified with $E\times Gr(l,r)$. Recall that the Hausdorff and upper box dimensions for the set $Gr(l,r)$ coincide (since the Grassmannian is a manifold); therefore, by a classical product formula (see for example [18], Corollary 7.4) $\dim_{H}P_{E}=\dim_{H}Gr(l,r)+\dim_{H}E=r(l-r)$. The set $X_{E}$ is the projection of $P_{E}$ on the second coordinate, so $\dim_{H}X_{E}\leq\dim_{H}P_{E}=r(l-r)$.
On the other hand, let us notice that the set $X_{\{e\}}$ of all $r$-dimensional subspaces containing a single point $e$ can be identified with $Gr(l,r)$ and thus has dimension r(lr)r(l-r). But $E$ is nonempty, hence $X_{E}$ contains a set of the form $X_{\{e\}}$ and therefore $\dim_{H}X_{E}\geq\dim_{H}X_{\{e\}}=r(l-r)$, which completes the proof. ∎
@@
Let $X_{E}$ be the set of $k$-dimensional linear subspaces of $\mathbb{R}^{n}$ intersecting with the set $E$ of $\Xi$-exceptional points for $\Theta^{\top}$. If $E$ is not empty, then $\dim_{H}X_{E}=(k-1)(n-k)$.
Corollary 3 shows that the linear subspaces $\mathcal{L}$ corresponding to $\Xi$-exceptional $k$-dimensional affine subspaces $\mathcal{A}$ form a $\leq(k-1)(n-k)$-dimensional subset on the grassmanian $Gr(n,k)$ for any $\Xi>0$. In particular, the set of $\Xi$-exceptional $k$-dimensional affine subspaces has measure zero.